OpenStudy (anonymous):

Please Help!! Use mathematical induction to prove the statement is true for all positive integers n, or show why it is false. 4*6+5*7+6*8+...+4n(4n+2)=(4(4n+1)(8n+7))/6

OpenStudy (anonymous):

@SolomonZelman

OpenStudy (anonymous):

@zzr0ck3r

OpenStudy (anonymous):

yes

OpenStudy (zzr0ck3r):

so you want to prove \(\sum_{i=1}^n4i(4i+1)=\frac{4(4n+1)(8n+7)}{6}\)

OpenStudy (zzr0ck3r):

ill write it out and upload one sec

OpenStudy (anonymous):

ok, thank you :)

OpenStudy (zzr0ck3r):

I dont think it does

OpenStudy (anonymous):

well it also says "or show why it is false"

OpenStudy (zzr0ck3r):

4*6+5*7+6*8 = 107 4(4*3+1)(8*3+7)/6 = 806/3

OpenStudy (zzr0ck3r):

so I looked for n = 3

OpenStudy (anonymous):

why does n=3

OpenStudy (zzr0ck3r):

I think you are confused on what we are doing you have a statement that says something is true for all n, I showed that it is not true for n = 3, thus it is NOT the case that it is true for all n

OpenStudy (anonymous):

so wouldn't it be 4*6+5*7+6*8+4(3)(4*3+2)=(4(4*3+1)(8*3+7))/6

OpenStudy (zzr0ck3r):

something is messed up for the sequence 4n(4n+2) we get 4*6, 8*10, 12*14 reread the question and make sure you posted it correctly

OpenStudy (anonymous):

Use mathematical induction to prove the statement is true for all positive integers n, or show why it is false. (4 points each.) 4 ⋅ 6 + 5 ⋅ 7 + 6 ⋅ 8 + ... + 4n( 4n + 2) = 4(4n+1)(8n+7)/6

OpenStudy (zzr0ck3r):

do you see why this is messed up?

OpenStudy (zzr0ck3r):

4*6+5*7+6*8 ... does not come from 4n(4n+2)

OpenStudy (anonymous):

why?

OpenStudy (zzr0ck3r):

let n=1 and you get 4*(4+2)=4*6 as this is good so far. let n = 2 and we get 4*2(4*2+1)=8*9 NOOOOO

OpenStudy (anonymous):

ohhh ok

OpenStudy (anonymous):

so is that what i would write?

OpenStudy (zzr0ck3r):

your sequence says we shuold get 4*6+5*7 not 4*6+8*10

OpenStudy (zzr0ck3r):

I would write that the question does not make sense.

OpenStudy (zzr0ck3r):

there is a typo

OpenStudy (anonymous):

can you help me with another one?

OpenStudy (zzr0ck3r):

I am not sure if we shuold consider the sequence 4*6+5*7+6*8 or 4*6+8*10+12*14...

OpenStudy (zzr0ck3r):

sure

OpenStudy (anonymous):

12 + 42 + 72 + ... + (3n - 2)^2 = n(6n^2-3n-1)/2

OpenStudy (anonymous):

@zzr0ck3r

OpenStudy (anonymous):

hey @wio can you help with this??

OpenStudy (zzr0ck3r):

lol I said I would

OpenStudy (anonymous):

no no i know you are but hes just sitting there...

OpenStudy (anonymous):

watching...

OpenStudy (zzr0ck3r):

again the question makes no sense....nothing squared is 12

OpenStudy (zzr0ck3r):

where are you getting these?

OpenStudy (anonymous):

my math assignment from flvs

OpenStudy (zzr0ck3r):

he sequence you showed shuold say 12+42+72+......+n^2-n

OpenStudy (zzr0ck3r):

starting with n = 4

OpenStudy (anonymous):

but it doesn't >.<

OpenStudy (zzr0ck3r):

@wio do you have any idea whats going on here?

OpenStudy (anonymous):

Is it the same question?

OpenStudy (zzr0ck3r):

they are not making sense

OpenStudy (zzr0ck3r):

12 + 42 + 72 + ... + (3n - 2)^2 = n(6n^2-3n-1)/2

OpenStudy (zzr0ck3r):

they are giving the wrong definition for the sequences (3n-2)^2 for this one.

OpenStudy (zzr0ck3r):

and something strange on the last one

OpenStudy (zzr0ck3r):

shuold say 12 + 42 + 72 + ... + n^2-n = starting at n=4

OpenStudy (anonymous):

@zzr0ck3r @wio

OpenStudy (aum):

For the first problem, the nth term they have given is wrong. The nth term should be: (n+3)(n+5) n = 1: 4*5 n = 2: 5*7 n = 3: 6*8

OpenStudy (aum):

It is not clear whether they actually want the sum of the series or just to prove whether the nth term is correct or not. The nth term is wrong on the first one because if you put n = 2 you get 8*10 which is not the second term in the series. The nth term on the second one is correct because if you put n = 1, you get 1^2, n = 2 gives 4^2, n = 3 gives 7^2.

OpenStudy (anonymous):

they just want to know if it is true or false and how i got to that conclusion but i didn't know how to go about it.

OpenStudy (anonymous):

@aum

OpenStudy (anonymous):

I see, so \(5\times 7\) is not a term in the sequence \(4n(4n+2)\).

OpenStudy (anonymous):

so what does 4n(4n+1)(8n+7)/6 have to do with anything

OpenStudy (anonymous):

Would the sequence be \((n+3)(n+5)\)

OpenStudy (anonymous):

i have no clue, i am so confused

OpenStudy (zzr0ck3r):

ok ninja this is how it would normally work 1+3+5+.....+(2n-1)=n^2 what that means is 1+3+5+7+9..... is a sequence that is given by the rule 2n-1 i.e. if we let n=1 we get 1, if we let n=2 we get 3, if we let n=3 we get 5. in the two series you gave us, the rule does not match the sequence before the rule

OpenStudy (zzr0ck3r):

@ninjasandtigers

OpenStudy (anonymous):

uh ok thank you. @zzr0ck3r